Find the equation of the cone whose vertex is at the origin and base is the circle x=a, y^2+z^2=b^2.
1
Expert's answer
2020-10-29T19:49:06-0400

The sphere S intersects with the plane П "\to" x = a gives the circle

"C \to y^2+z^2=b^2"

"S \to ||p-p_0||=r"

with

"r = \frac{b}{a}\sqrt{b^2+a^2}"

"p = (x,y,z), p_0 = (x_0,y_0,z_0)"

"x_0 = a+\sqrt{r^2-b^2}, y_0=z_0=0"

Consider the line "L\to p=\lambda \vec{v}" which passes by (0,0,0)

intersections S and L:

"||\lambda\vec{v} - p_0||=r"

"\lambda^2||\vec{v}||^2-2\lambda\langle\vec{v}, p_0\rangle+||p_0||^2=r^2"

"\lambda=\langle\vec{v}, p_0\rangle \pm \sqrt{\langle\vec{v}, p_0\rangle^2-||\vec{v}||^2(||p_0||^2-r^2)}"

the line L must be tangent to S hence

"\langle\vec{v}, p_0\rangle^2-||\vec{v}||^2(||p_0||^2-r^2)=0"

This is the locus defining the cone surface.

Putting "\vec{v}=(x,y,z)" and applying the values for "p_0,r" we get

"b^4x^2-a^4(y^2+z^2)-2a^2b^2(y^2+z^2)+a^2b^2(x^2+y^2+z^2)=0"